OpenStudy (anonymous):

What is the derivative of y=xsinx+cosx? Is it simply y'=1cosx-sinx because derivative of x=1, derivative of sinx=cosx and derivative of cosx=-sinx?

OpenStudy (anonymous):

\[y' = [x(\cos x) + \sin x] - \sin x = x(\cos x)\]You just use the product rule on that first term to get the value within the brackets.

OpenStudy (anonymous):

If y = uv then y' = u'v + uv'

OpenStudy (amistre64):

only if you are taking the derivative "with respect to x" is x' = 1

OpenStudy (anonymous):

All good now, @schi1 ?

OpenStudy (anonymous):

still not sure where the extra -sinx came from

OpenStudy (anonymous):

the derivative of the cos is the negative sine

OpenStudy (anonymous):

the derivative of sinx should equal cosx and the derivative of cosx should equal -sinx shouldn't it?

OpenStudy (anonymous):

so the derivative of y=xsinx+cosx should just be cosx-sinx isn't it?

OpenStudy (amistre64):

y = xf+g y' = (xf)' + (g)' y' = x'f + xf' + g' , if this is wrt.x, x'=1 y' = f + xf' + g'

OpenStudy (anonymous):

Yes, that's exactly right, @schi . In that original function, you have 2 terms. The first one shows the derivative in the brackets. You still have to do the derivative of the second term.

OpenStudy (anonymous):

ohh ok gotcha I missed a rule

OpenStudy (anonymous):

In your problem, the way it is written, that "x" gets multiplied only to sin x indicating the derivation that @amistre64 so very well pointed out. If that is how the problem is indeed intended. If, however you want "x" multiplied to the sum of sinx and cos x, you have to put that in parentheses.

OpenStudy (anonymous):

no x only multiplies sinx

OpenStudy (anonymous):

ok, that's good. Then that first post of mine gives the actual derivative.

OpenStudy (anonymous):

So, to sum up, you get the derivative of x sin x -> x cos x + sin x And then you get the derivative of cos x -> - sin x And then you add those. The sin x and the "- sin x" will cancel.

OpenStudy (anonymous):

You should be all set now, @schi1 . Are you ok with the answer now?

OpenStudy (anonymous):

yes thank you